Average age of 4 sons of a family is 12 years. Average age of sons together with their parents is 28 years. If father is older than mother by 6 years . Find the age of mother?
Answer by lwsshak3(11628)

Average age of 4 sons of a family is 12 years. Average age of sons together with their parents is 28 years. If father is older than mother by 6 years . Find the age of mother?
**
let x=age of mother
x+6=age of father
..
Total age of 4 sons/4=12
Total age of 4 sons=48
..
(Total age of 4 sons+father's age+mother's age)/6=28
(48+x+6+x)/6=28
48+x+6+x=168
2x=168-54=114
x=114/2
x=57
ans:
age of mother=57